# AdaNCA: Neural Cellular Automata As Adaptors For More Robust Vision Transformer

### v4

@@ Abstract
Vision Transformers (ViTs) have demonstrated remarkable performance in image classification tasks, particularly when equipped with local information via region attention or convolutions. While such architectures improve the feature aggregation from different granularities, they often fail to contribute to the robustness of the networks. Neural Cellular Automata (NCA) enables the modeling of global cell representations through local interactions, with its training strategies and architecture design conferring strong generalization ability and robustness against noisy input. In this paper, we propose Adaptor Neural Cellular Automata (AdaNCA) for Vision Transformer that uses NCA as plug-in-play adaptors between ViT layers, enhancing ViT’s performance and robustness against adversarial samples as well as out-of-distribution inputs. To overcome the large computational overhead of standard NCAs, we propose Dynamic Interaction for more efficient interaction learning. Furthermore, we develop an algorithm for identifying the most effective insertion points for AdaNCA based on our analysis of AdaNCA placement and robustness improvement. With less than a 3% increase in parameters, AdaNCA contributes to more than 10% absolute improvement in accuracy under adversarial attacks on the ImageNet1K benchmark. Moreover, we demonstrate with extensive evaluations across 8 robustness benchmarks and 4 ViT architectures that AdaNCA, as a plug-in-play module, consistently improves the robustness of ViTs.
@@ 1 Introduction
Vision Transformers (ViTs) exhibit impressive performance in image classification through globally modeling token interactions via self-attention mechanisms [11, 14, 76]. Recent works show that integrating local information into ViTs, e.g., using region attention  [25, 37, 38, 53, 64, 69, 72] or convolution [8, 10, 16, 21, 22, 36, 41, 51, 67, 70, 74], further enhances the ViT’s capabilities in image classification. While advanced local structures contribute to better captures of local information, the robustness of ViTs has not increased. They remain vulnerable to noisy input such as adversarial samples [5, 9, 13, 40, 41] and out-of-distribution (OOD) inputs [17, 18, 27, 29, 66].
@@
Recently, Neural Cellular Automata (NCA) has been proposed as a lightweight architecture for modeling local cell interactions [43], where cells are represented by 1D vectors. Similar to the idea of token interactions in ViTs, cells in NCA interact with each other by alternating between a convolution-based Interaction stage and an MLP-based Update stage [45, 48] to perform downstream tasks. The critical difference, however, is that cell interactions in NCA evolve over time by recurrent application of the two stages, while ViT computes the token interaction in a single step per layer. During this process, cells dynamically modulate their representations based on the interactions with their neighbors and gradually enlarge their receptive fields. Unlike commonly used convolutional neural networks, NCA maintains resolution during neighborhood expansion, and the parameter sharing across the loop further regularizes the learning process, preventing NCA from overfitting and enhancing its generalization ability [43, 45]. Moreover, NCA training involves various kinds of stochasticity [44], enabling the models to generalize to input variability and adapt to unpredictable perturbations. It is the modulation of local information and stochasticity during training that make NCA robust against noisy input [48, 47, 54, 61].
@@
Figure 1: The accuracy under adversarial attacks (APGD-DLR [9]) v.s. corruption error on out-of-distribution input (ImageNet-C [27]) of various ViT models [16, 22, 37, 41]. AdaNCA improves the robustness of different ViTs against both adversarial attacks and OOD input. $\star$: the same model architecture but with more layers.
@@
However, the original NCA has substantial computational overhead when operating in high-dimensional space, a common scenario in ViTs. This poses a non-trivial challenge when integrating NCA into ViTs. To reduce the dimensionality of interaction results and thus lower the computational cost, we propose Dynamic Interaction to replace the standard Interaction stage. In addition, tokens dynamically modify the interaction strategy based on the observation of their environment. This adaptation to the variability of the environment contributes to the model robustness. Our modified Adaptor NCA (AdaNCA), as a plug-in-play module, improves ViTs performances as illustrated in Figure 1. Adding AdaNCA to different ViT architectures consistently improves their robustness to both adversarial attacks and OOD input. AdaNCA also improves clean accuracy.
@@
Motivated by empirical observations of the positive relationship between network redundancy and model robustness [27], we develop a dynamic programming algorithm to compute the most effective insert position for AdaNCA within a ViT, based on our proposed quantification of network redundancy. Our method results in consistent improvements across 8 robustness benchmarks and 4 different baseline ViT architectures. Critically, we demonstrate that the improvements do not originate from the increase in parameters and FLOPs, but are attributed to AdaNCA. Our contributions are as follows:
@@
  • We propose AdaNCA, which integrates Neural Cellular Automata into ViTs’ middle layers as lightweight Adaptors to enhance the robustness of ViTs against adversarial attacks and OOD inputs in image classification. With less than 3% more parameters, AdaNCA-extended ViTs can achieve 10% higher accuracy under certain adversarial attacks.
@@
  • We introduce Dynamic Interaction to replace the Interaction stage in standard NCA, enhancing model robustness and efficiency in terms of parameters and computation.
@@
  • We propose a method to determine the most effective insert positions of AdaNCA for maximum robustness improvement.
@@
Figure 2: Method overview. (a) Neural Cellular Automata (NCA) can be inserted into Vision Transformers (ViTs) as Adaptors to improve model performance and robustness, hence termed AdaNCA. The improvement is maximized when AdaNCA is inserted between two layer sets that each consists of similar layers. (b) A single step of AdaNCA. Our proposed Dynamic Interaction contributes to model efficiency as well as robustness. (c) The robustness improvement brought by AdaNCA is highly correlated with the corresponding network redundancy quantification of the insert position. It supports that AdaNCA should be placed between two sets of redundant layers.
@@ 2 Related Works
2.1 Local structure in Vision Transformer
Since the proposal of Vision Transformer (ViT) [14], a series of works have introduced local structures into ViTs to enhance their performance [8, 10, 16, 21, 22, 36, 41, 51, 67, 69, 70, 72, 74, 75]. Here, we mention one of the earliest local structure modifications and the ones relevant to our work. Stand Alone Self Attention (SASA), as introduced by Ramachandran et al. [53], utilizes sliding window self-attention in ViTs. Following this, Liu et al. [37] develop a non-sliding window attention mechanism that partitions feature maps and computes self-attention both within and between these partitions, termed Shifted Window (Swin) attention. Another method to model local information is convolution. d’Ascoli et al. [16] introduce soft local inductive bias using gated positional self-attention, fusing self-attention and convolution. Despite these advancements in better modeling of local information, few methods contribute to a more robust ViT architecture [41], leaving the models to behave subpar when encountering slightly noisy inputs or distribution shifts.
2.2 Robust architecture in Vision Transformer
Researchers have developed various architectural changes to build more robust ViTs against adversarial attacks, such as FGSM [60] or PGD [40], as well as out-of-distribution (OOD) inputs, such as image corruption [27]. Zhou et al. [77] propose Full Attention Networks to boost the robustness of ViTs against OOD images. Mao et al. [41] first systematically analyze the relationship between different components in a ViT, drawing a positive relationship between convolutional patch embedding as well as convolution in MLP and the robustness of ViTs against adversarial samples and OOD data. Extending [41] and [77], Guo et al. [22] propose input-dependent average pooling to adaptively select different aggregation neighborhoods for different tokens, achieving the state-of-the-art robust ViTs in OOD generalization. Different interpretations of the self-attention operation can also lead to more robust architectures [23, 57]. However, those methods that introduce additional architectures [22, 23, 57, 77] are either implemented on ViTs with limited size or focus on non-adversarial robustness. On the contrary, our method introduces NCA as lightweight plug-in-play adaptors into base-level ViTs, enhancing their clean accuracy and robustness against both adversarial samples and OOD inputs.
@@ 2.3 Neural Cellular Automata
Mordvintsev et al. [43] propose NCA to mimic the biological cell interactions and model morphogenesis. Following this, several works have applied NCA in computer vision, including texture synthesis [42, 45, 48, 47, 49], image generation [32, 46, 50, 61], and image segmentation [31, 56]. Randazzo et al. [54] propose applying NCA for modeling collective intelligence on image classification tasks, although the limitation to binary images restricts its practical application. Tesfaldet et al. [61] first establish the connection between ViT and NCA via recurrent local attention. However, their application is limited to image impainting on small datasets such as MNIST [1] and CIFAR10 [35]. Our work first applies NCA in image classification on ImageNet1K with base-level ViT models. Moreover, we propose the new Dynamic Interaction for efficient cell interaction modeling, reducing the computational overhead and enhancing the model performance.
@@ 3 Method
The overview of our method is shown in Figure 2. In this section, we first review the NCA model and ViT architecture in Section 3.1 and establish the connection between NCA and ViT in terms of token interaction modeling. We then present the design of our AdaNCA in Section 3.2. We insert AdaNCA into the middle layers of ViT to improve its robustness, and we introduce the relationship between the insert position of AdaNCA and the relative improvements of model robustness in Section 3.3, which leads to the algorithm for deciding the most effective placement for AdaNCA.
@@ 3.1 Preliminaries
Vision Transformer
Vision Transformers (ViTs) operate on token maps $\mathbf{X}\in\mathbb{R}^{N\times C}$, where the number of tokens is $N$ and the each token is represented by a $C$-dimensional vector. ViTs learn the interaction between these tokens via self-attention [63] and computes the interaction result $\mathbf{X}_{attn}$, as described in Equation $1$We ignore the residual update scheme for simplicity.
$\mathbf{X}_{attn}=\sigma\left(\frac{\mathbf{Q}\mathbf{K}^{\top}}{\sqrt{D}}\right)\mathbf{V}.$ (1)
$\mathbf{Q},\mathbf{K},\mathbf{V}$ stand for query, key, and value, respectively. They are deduced from different linear projections of the input, i.e., $\mathbf{Q}=\mathbf{X}\mathbf{W}_{\mathbf{Q}},\mathbf{K}=\mathbf{X}\mathbf{W}_{\mathbf{K}},\mathbf{V}=\mathbf{X}\mathbf{W}_{\mathbf{V}}$, where $\mathbf{W}_{\mathbf{Q}},\mathbf{W}_{\mathbf{K}},\mathbf{W}_{\mathbf{V}}\in\mathbb{R}^{C\times D}$. $D$ is the hidden dimensionality in self-attention. $\sigma$ is Softmax. After self-attention, tokens are fed into a Multilayer Perceptron (MLP) to get the updated representations $\mathbf{X}_{out}$:
@@
$f$ is the MLP and $\theta$ stands for its parameters. The self-attention and MLP form a single ViT block, and one can build a ViT model via stacking ViT blocks.
@@
NCA aims at modeling cell interactions. In the 2D domain, cells live on a 2D grid with size $H\times W$. Each cell is represented by a vector with dimensionality $C$. All cells collectively define the cell states $\mathbf{S}\in\mathbb{R}^{H\times W\times C}$. In a single step of NCA, cells first interact with their neighborhoods for information exchange in the Interaction stage [48] to generate the interaction output $\mathbf{S}_{\mathcal{I}}$, typically instantiated via depth-wise convolutions [43, 45, 47]:
@@
$\mathcal{C}_{i}$ is the $i$th convolutional kernel, and $\mathcal{M}$ denotes the total number of kernels. The kernels can either be fixed [48, 47] or learnable [43, 45]. The results of all kernels are concatenated channel-wise in the $\oplus$ operation. $\mathbf{S}_{\mathcal{I}}$ is then passed to an MLP in the Update stage [48], where cells get the updated states $\mathbf{S}_{out}$:
@@
$f$ is the MLP and $\theta$ stands for its parameters. Typically, NCA uses the simplest MLP with two linear layers and one activation between them. $\mathbf{M}\in\mathbb{R}^{H\times W\times C}$, sampled from $Bernoulli(p)$, is a random binary mask to introduce stochasticity in NCA, ensuring asynchronicity during the cell updates [44]. $\odot$ is point-wise multiplication. NCA learns an underlying dynamic governing the cell behaviors [49], as depicted by the stochastic differential equation (SDE) in Equation 5:
@@
NCA outputs the cell states at time $t=T$ to accomplish downstream tasks. $T$ is a randomly-chosen integer in a certain range.
@@ 3.1.1 Connecting NCA and ViT
Both NCA and ViT learn interactions between a set of elements, i.e., tokens in ViT and cells in NCA. From now on, we refer to a cell in NCA as a token, aligning it with the concept in ViTs. The asynchronicity [44] introduced by the random mask $\mathbf{M}$ can be regarded as a cell-wise stochastic depth [30], a more fine-grained version of the sample-wise stochastic depth. In previous NCA works, stochasticity is maintained during testing [45]. Such a scheme is problematic in discriminative tasks since reliable numerical outputs are essential. To this end, we adopt the strategy in dropout-like techniques [58, 30] that compensates activation values during training. Given $\mathbf{M}\sim Bernoulli(p)$, the evolution of NCA in AdaNCA is defined as:
@@
Furthermore, NCA typically queries the cell states at a random time step $T$, resulting in random update steps for all cells. Such randomness ensures the stability of NCA across various time steps [43]. Hence, the trained model can effectively handle the variability and unpredictability of the input, thus being robust against noisy input. Our ablation studies in Section 4.3 demonstrate the effectiveness of these two strategies in enhancing the model performance and robustness.

### v5

@@ Abstract
Vision Transformers (ViTs) demonstrate remarkable performance in image classification through visual-token interaction learning, particularly when equipped with local information via region attention or convolutions. Although such architectures improve the feature aggregation from different granularities, they often fail to contribute to the robustness of the networks. Neural Cellular Automata (NCA) enables the modeling of global visual-token representations through local interactions, as its training strategies and architecture design confer strong generalization ability and robustness against noisy input. In this paper, we propose Adaptor Neural Cellular Automata (AdaNCA) for Vision Transformers that uses NCA as plug-and-play adaptors between ViT layers, thus enhancing ViT’s performance and robustness against adversarial samples as well as out-of-distribution inputs. To overcome the large computational overhead of standard NCAs, we propose Dynamic Interaction for more efficient interaction learning. Using our analysis of AdaNCA placement and robustness improvement, we also develop an algorithm for identifying the most effective insertion points for AdaNCA. With less than a 3% increase in parameters, AdaNCA contributes to more than 10% of absolute improvement in accuracy under adversarial attacks on the ImageNet1K benchmark. Moreover, we demonstrate with extensive evaluations across eight robustness benchmarks and four ViT architectures that AdaNCA, as a plug-and-play module, consistently improves the robustness of ViTs.
@@ 1 Introduction
Vision Transformers (ViTs) exhibit impressive performance in image classification, through globally modeling token interactions via self-attention mechanisms [11, 14, 77]. Recent works show that integrating local information into ViTs, e.g., using region attention  [26, 38, 39, 54, 65, 70, 73] or convolution [8, 10, 16, 22, 23, 37, 42, 52, 68, 71, 75], further enhances the ViT’s capabilities in image classification. Although advanced local structures contribute to better captures of local information, the robustness of ViTs has not increased. They remain vulnerable to noisy input such as adversarial samples [5, 9, 13, 41, 42] and out-of-distribution (OOD) inputs [17, 18, 28, 30, 67].
@@
Recently, Neural Cellular Automata (NCA) was proposed as a lightweight architecture for modeling local cell interactions [44], where cells are represented by 1D vectors. To perform downstream tasks, similarly to the idea of token interactions in ViTs, cells in NCA interact with each other by alternating between a convolution-based Interaction stage and an MLP-based Update stage [46, 49]. The critical difference, however, is that cell interactions in NCA evolve over time by recurrent application of the two stages, whereas ViT computes the token interaction in a single step per layer. During this process, cells dynamically modulate their representations, based on the interactions with their neighbors, and they gradually enlarge their receptive fields. Unlike commonly used convolutional neural networks, NCA maintains resolution during neighborhood expansion. The recurrent update scheme enables the cells to explore various states, thus preventing NCA from overfitting and enhancing its generalization ability [44, 46]. NCA training involves various kinds of stochasticity [45], which enables the models to generalize to input variability and adapt to unpredictable perturbations. It is the modulation of local information and stochasticity during training that make NCA robust against noisy input [49, 48, 55, 62].
@@
Figure 1: The accuracy under adversarial attacks (APGD-DLR [9]) versus corruption error on out-of-distribution input (ImageNet-C [28]) of various ViT models [16, 23, 38, 42]. AdaNCA improves the robustness of different ViTs against both adversarial attacks and OOD input. $\star$: the same model architecture but with more layers.
@@
However, the original NCA has substantial computational overhead when operating in high-dimensional space, which is a common scenario in ViTs. This poses a non-trivial challenge when integrating NCA into ViTs. To reduce the dimensionality of interaction results, hence to lower the computational cost, we propose Dynamic Interaction to replace the standard Interaction stage. In this stage, tokens dynamically modify the interaction strategy, based on the observation of their environment. This adaptation to the variability of the environment contributes to the model robustness. Our modified Adaptor NCA (AdaNCA), as a plug-and-play module, improves ViTs performances, as illustrated in Figure 1. Adding AdaNCA to different ViT architectures consistently improves their robustness to both adversarial attacks and OOD input. AdaNCA also improves clean accuracy.
@@
Motivated by empirical observations of the positive relationship between network redundancy and model robustness [28], we develop a dynamic programming algorithm for computing the most effective insert position for AdaNCA within a ViT, based on our proposed quantification of network redundancy. Our method results in consistent improvements across eight robustness benchmarks and four different baseline ViT architectures. Critically, we demonstrate that the improvements do not originate from the increase in parameters and FLOPS but are attributed to AdaNCA. Our contributions are as follows:
@@
  • We propose AdaNCA: It integrates Neural Cellular Automata into ViTs’ middle layers as lightweight Adaptors for the robustness enhancement of ViTs against adversarial attacks and OOD inputs in image classification. With less than 3% more parameters, AdaNCA-extended ViTs can, under certain adversarial attacks, achieve 10% higher accuracy.
@@
  • We introduce Dynamic Interaction to replace the Interaction stage in standard NCA, thus enhancing model robustness and efficiency in terms of parameters and computation.
@@
  • We propose a method for determining the most effective insert positions of AdaNCA, for maximum robustness improvement.
@@
Figure 2: Method overview. (a) To improve model performance and robustness, Neural Cellular Automata (NCA) can be inserted into Vision Transformers (ViTs) as Adaptors, hence termed AdaNCA. The details of AdaNCA are presented in Section 3.2. The improvement is maximized when AdaNCA is inserted between two layer sets that each consists of similar layers. (b) The robustness improvement brought by AdaNCA is highly correlated with the corresponding network redundancy quantification of the insert position introduced in Section 3.3. This supports the idea that AdaNCA should be placed between two sets of redundant layers.
@@ 2 Related Works
2.1 Local structure in Vision Transformers
Since the proposal of Vision Transformer (ViT) [14], a series of works have introduced local structures into ViTs to enhance their performance [8, 10, 16, 22, 23, 37, 42, 52, 62, 68, 70, 71, 73, 75, 76]. Here, we mention one of the earliest local structure modifications and those relevant to our work. Stand-Alone Self-Attention (SASA), as introduced by Ramachandran et al. [54], utilizes sliding window self-attention in ViTs. Following this, Liu et al. [38] develop a non-sliding window attention mechanism that partitions feature maps and computes self-attention, both within and between these partitions; it is termed Shifted Window (Swin) attention. Another method for modeling local information is convolution. D’Ascoli et al. [16] introduce a soft local-inductive bias by using gated positional self-attention, thus fusing self-attention and convolution. Despite these advancements in better modeling of local information, few methods lead to a more robust ViT architecture [42]. This leaves the models to behave subpar when encountering slightly noisy inputs or distribution shifts.
2.2 Robust architecture in Vision Transformers
Researchers have developed various architectural changes for building more robust ViTs against adversarial attacks, such as FGSM [61] or PGD [41], as well as out-of-distribution (OOD) inputs, such as image corruption [28]. Zhou et al. [78] propose Full Attention Networks to boost the robustness of ViTs against OOD images. Mao et al. [42] first systematically analyze the relationship between different components in a ViT, drawing a positive relationship between convolutional components and the robustness of ViTs against adversarial samples and OOD data. By extending [42] and [78], Guo et al. [23] propose input-dependent average pooling in order to adaptively select different aggregation neighborhoods for different tokens, thus achieving the state-of-the-art robust ViTs in OOD generalization. Different interpretations of the self-attention operation can also lead to more robust architectures [24, 58]. However, the methods that introduce additional architectures [23, 24, 58, 78] are either implemented on ViTs with limited size or focus on non-adversarial robustness. On the contrary, our method introduces NCA as lightweight plug-and-play adaptors into base-level ViTs, thus enhancing their clean accuracy and robustness against both adversarial samples and OOD inputs.
@@ 2.3 Neural Cellular Automata
Gilpin [20] demonstrates that CA can be represented by convolutional neural networks. By extending [20], Mordvintsev et al. [44] propose NCA in order to mimic the biological cell interactions and model morphogenesis. Following this idea, several works apply NCA in computer vision, including texture synthesis [43, 46, 49, 48, 50], image generation [33, 47, 51, 62], and image segmentation [32, 57]. Randazzo et al. [55] propose applying NCA for modeling collective intelligence on image classification tasks, but the limitation to binary images restricts its practical application. Tesfaldet et al. [62] first establish the connection between ViT and NCA via recurrent local attention. It leads to a more robust model for handling image corruptions in image inpainting tasks. However, their application is limited to image impainting on small datasets such as MNIST [1] and CIFAR10 [36]. The NCA in [62] attempts to emulate a ViT whereas our approach distinguishes itself by not doing so. We first applies NCA in image classification on ImageNet1K with base-level ViT models. Moreover, we propose the new Dynamic Interaction for efficient cell interaction modeling, reducing the computational overhead, and for enhancing the model performance.
@@ 3 Method
The overview of our method is shown in Figure 2. In this section, we first review the NCA model and ViT architecture. In Section 3.1, we establish the connection between NCA and ViT, in terms of token interaction modeling. We then present the design of our AdaNCA in Section 3.2. We insert AdaNCA into the middle layers of ViT to improve its robustness. We introduce the relationship between the insert position of AdaNCA and the relative improvements of model robustness in Section 3.3. This relationship leads to the algorithm for deciding the most effective placement for AdaNCA.
@@ 3.1 Preliminaries
Vision Transformers
Vision Transformers (ViTs) operate on token maps $\mathbf{X}\in\mathbb{R}^{N\times C}$, where the number of tokens is $N$ and each token is represented by a $C$-dimensional vector. ViTs learn the interaction between these tokens via self-attention [64] and compute the interaction result $\mathbf{X}_{attn}$, as described in Equation 1.
$\mathbf{X}_{attn}=\sigma\left(\frac{\mathbf{Q}\mathbf{K}^{\top}}{\sqrt{C}}\right)\mathbf{V}.$ (1)
$\mathbf{Q},\mathbf{K},\mathbf{V}$ stand for query, key, and value, respectively. They are deduced from different linear projections of the input, i.e., $\mathbf{Q}=\mathbf{X}\mathbf{W}_{\mathbf{Q}},\mathbf{K}=\mathbf{X}\mathbf{W}_{\mathbf{K}},\mathbf{V}=\mathbf{X}\mathbf{W}_{\mathbf{V}}$, where $\mathbf{W}_{\mathbf{Q}},\mathbf{W}_{\mathbf{K}},\mathbf{W}_{\mathbf{V}}\in\mathbb{R}^{C\times D}$. $D$ is the hidden dimensionality in self-attention. $\sigma$ is Softmax. After self-attention, tokens are fed into a Multilayer Perceptron (MLP) to obtain the updated representations $\mathbf{X}_{out}$:
@@
$f$ is the MLP and $\theta$ stands for its parameters. The self-attention and MLP form a single ViT block, and a ViT model can be built via stacking ViT blocks.
@@
NCA aims at modeling cell interactions. In the 2D domain, cells live on a 2D grid with size $H\times W$. Each cell is represented by a vector with dimensionality $C$. All cells collectively define the cell states $\mathbf{S}\in\mathbb{R}^{H\times W\times C}$. In a single step of NCA, to generate the interaction output $\mathbf{S}_{\mathcal{I}}$, cells first interact with their neighborhoods for an information exchange in the Interaction stage [49]; the interaction is typically instantiated via depth-wise convolutions [44, 46, 48]:
@@
[…]
@@
$\mathbf{S}_{out}$ is then used to update the cell states in a residual scheme. $f$ is the MLP, and $\theta$ stands for its parameters. Typically, NCA uses the simplest MLP, with two linear layers and one activation between them. $\mathbf{M}\in\mathbb{R}^{H\times W\times C}$, sampled from $Bernoulli(p)$, is a random binary mask to introduce stochasticity in NCA; it ensures asynchronicity during the cell updates [45]. $\odot$ is point-wise multiplication. NCA learns an underlying dynamic that governs the cell behaviors [50], as depicted by the stochastic differential equation (SDE) in Equation 5:
@@
After $t=T$ steps, the cell states $S^{T}$ is extracted to accomplish certain downstream tasks. The traditional NCA involves several other specific designs though, in our case, it is impractical to adapt them. We provide a discussion on this topic in Appendix E.
@@ 3.1.1 Connecting NCA and ViT
Both NCA and ViT learn interactions between a set of elements, i.e., tokens in ViT and cells in NCA. Hereafter, we refer to a cell in NCA as a token, aligning it with the concept in ViTs. The asynchronicity [45] introduced by the random mask $\mathbf{M}$ can be regarded as a cell-wise stochastic depth [31], a more fine-grained version of the sample-wise stochastic depth. In previous NCA works, stochasticity is maintained during testing [46]. Such a scheme is problematic in our case because (1) test-time stochasticity produces obfuscated gradients [3], leading to the circumvention of adversarial attacks, and (2) the model can output different results given the same inputs. To this end, we adopt the strategy in dropout-like techniques [31, 59], which compensates activation values during training. Given $\mathbf{M}\sim Bernoulli(p)$, the evolution of NCA in AdaNCA is defined as:
@@
We discuss the necessity of such a scheme in Section D.1 in the Appendix. Furthermore, NCA typically outputs the cell states at a random time step $T$, resulting in random update steps for all cells. Such randomness ensures the stability of NCA across various time steps [44]. Finally, the recurrent steps of NCA during a single training epoch enable the exploration of a wide range of cell states. In the early stage of training, the model is not adequately trained hence serves as a source of noise to itself through the recurrence. With all these components, the trained model can effectively handle the variability and unpredictability of the input thus be robust against noisy input. Our ablation studies in Section 4.3 demonstrate the effectiveness of these strategies in enhancing the model performance.